OpenStudy (anonymous):

If d varies inversely as t, and d=20 when t=2, what is the value of t when d=-5? Please show step by step:)

OpenStudy (wolf1728):

if d varies inversely, then t = (1/d) * x when d = 20 t =2 therefore x = 40 So, t = (x/d) Checking the formula 2 = (40/20) - formula is okay So if d=-5 then t = (40/-5) t = -8